OpenStudy (anonymous):

Multiply: (m + 7) (m - 7) the 7 crosses out right? so is the answer m^2-49?

OpenStudy (diyadiya):

m^2 - 49 is right!

OpenStudy (diyadiya):

But what 7 crossing out ? o.O

Parth (parthkohli):

Another tip. There's an identity - \((a + b)(a - b) = a^2 - b^2\)

OpenStudy (anonymous):

(a + b)(a - b) = \[a^2 - b^2\] Therefore, answer will be :\[m^2 - 49\]...